# Controlling the Exchange Field of Surface Spin Impurities via DC Voltages

### v2

@@ Abstract
Recent advances in scanning probe microscopy have demonstrated the quantum-coherent manipulation of individual spins on surfaces using oscillating AC voltages. At the same time, the role of the static voltage in such control schemes remains unclear. Here, we show how the static contribution can be utilized to control the resonance of electronic transitions via an effective exchange field model based on a non-equilibrium transport formalism. We demonstrate that the magnitude and direction of the exchange field can be tuned and turned off by changing the applied bias voltage. Furthermore, our model can accurately reproduce the observed resonance frequency shifts at different magnetic field geometries for two different spin $S=1/2$ adsorbates. This work enables fast and precise control of the quantum states of surface spins at the atomic scale which will aid in developing spin-based quantum technologies.
@@
Introduction– The manipulation and control of individual quantum spins on surfaces is critical for advancing quantum information processing and spintronic devices at the nanoscale [1, 2]. While oscillating magnetic fields delivered via strip lines have traditionally been used to control the quantum states of spins, electric fields offer significant advantages, including faster switching speeds and more localized application through local gate electrodes [3, 4, 5, 6, 7, 8, 9]. This is particularly crucial for quantum computing applications, where rapid gate operations are essential to mitigate decoherence effects. Furthermore, electric field control could lead to more energy-efficient spintronic devices, benefiting low-power, high-density information storage and processing technologies [4, 10]. Despite these potential advantages, the impact of local electric fields on the quantum properties of surface spins remains elusive. To investigate this, the recently developed combination of electron spin resonance (ESR) and scanning tunneling microscopy (STM) can be employed. ESR-STM allows for probing and manipulating single atomic and molecular spins with atomic scale precision and purely by electric fields [11, 12, 13, 14, 15], making it an ideal tool for studying the interplay between electric fields and spin dynamics on surfaces.
In this Letter, we focus on the frequency shift of the resonance of dipole-allowed ESR transitions of single-spin impurities in an STM junction formed by a magnetic tip, a magnetic adsorbate, and a substrate. When the tip is sufficiently close to the impurity, the interaction between the spin and the tip is dominated by the exchange interaction $B_{\mathrm{exch}}$ [12, 16], which leads to a shift in the resonance frequency that can be controlled by the tip-atom distance [12, 17]. Similar resonance frequency shifts have been induced purely by varying the DC bias at fixed tip-atom distance. This effect was originally attributed to $g$-factor modulation [18, 19], however the underlying physics remained elusive and may be related to the driving mechanism of all-electric ESR in STM [20].
We found that all experimental observations can be described using a transport formalism and the concept of a bias-dependent exchange field. This field originates from spin-dependent virtual quantum charge fluctuations in the transport junction, leading to spin splitting as observed in quantum dots coupled to ferromagnetic leads [21, 22, 23]. This formalism predicts a peculiar bias dependence of the exchange field that leads to a nonlinear resonance shift, an inversion of the shift direction when the tip polarization changes sign, a dependence of the shift on the angle between tip polarization and adsorbate spin, and a vanishing shift at non-zero DC bias. Specifically, we will show that a single orbital Anderson impurity model (SAIM) with time-dependent couplings to a magnetic tip [20] encapsulates all the features described above and accurately captures experimental data obtained from single Ti atoms and iron (II) phthalocyanine molecules. Our work paves the way for ultra-fast, atomically localized control of spin transitions via DC bias in transport-based quantum devices.
@@
Experimental Results– In our experiment, we utilized two adsorbates, FePc molecules and Ti atoms (Fig. 1(a)), which have electron spin $S=1/2$ when adsorbed on a bilayer of MgO grown on top of an Ag(100) substrate [13, 12]. The external magnetic field in this case was applied along out-of-plane direction ($B_{\mathrm{ext}}=B_{z}$) which, for a $S=1/2$, defines the spin quantization axis since the external magnetic field is much larger than the exchange field (see SM, Sec. 4). Consequently, the resonance frequency is determined by
$hf_{0}=g\mu_{B}(B_{\text{ext}}+B_{\text{exch}}^{\parallel}),$ (1)
@@
with $h$ the Planck constant, $g$ the $g$-factor and $\mu_{B}$ the Bohr magneton. We performed ESR measurements on individual spins by adding a radio-frequency AC voltage ($V_{\text{RF}}$) in addition to $V_{\text{DC}}$ to the STM tip, as illustrated in Fig. 1(b). We follow the convention that $V_{\text{DC}}<0$ corresponds to probing the filled states of the sample.
@@
Figure 1: ESR frequency ($f_{0}$) shift of an individual FePc molecule as a function of the DC bias. (a) STM image of the Fe, Ti atoms and the FePc molecule co-adsorbed on the MgO/Ag(100) surface. Scanning parameters: $V_{\text{DC}}=200$ mV, $I_{\text{set}}=20$ pA. (b) Schematic diagram of the ESR set-up at the tunneling junction, integrated into a plate capacitor model. $\vec{P}_{\mathrm{T}}$ represents the spin-polarization of the tip, forming an angle $\theta$ with the out-of-plane direction. (c) Experimentally measured magnetic moments and resonance frequencies $f_{0}$ (d) at different $V_{\text{DC}}$. The red line gives the weighted average value. The tip height was fixed with respect to the tunneling parameters of $V_{\text{DC}}=100$ mV, $I_{\text{set}}=20$ pA. ESR conditions: $V_{\text{RF}}=20$ mV, $B_{z}=560$ mT, $T=1.8$ K. (e) DFT-calculated $g_{z}$ for a FePc molecule as a function of an applied out-of-plane DC electric field. (f) Simulated $f_{0}$-shift as a function of $V_{\text{DC}}$ based on the $g_{z}$ modulation in (e).
@@
One possible explanation of the $V_{\mathrm{DC}}$ dependence of the resonance frequency is a bias dependence of $g$ in Eq. (1), i.e. $g=g(V_{\mathrm{DC}})$ [18, 19]. To estimate the $g$-factor variation in FePc we measured the magnetic moment $\mu$ at varying external fields. Here, $\mu$ remains largely constant across different $V_{\text{DC}}$ values (Fig. 1(c)) indicating that the $g$-factor does not significantly change within the bias range. Simultaneously, we observed a significant monotonic decrease in resonance frequency $f_{0}$ with increasing $|V_{\text{DC}}|$ at fixed tip height (Fig. 1(d)), where $B_{\text{exch}}^{\parallel}$ should remain constant, ruling out the $g$-factor explanation completely  [^1] We note that a precise determination of the $g$-factor from ESR-STM measurements is challenging due to competing influence of external and tip fields [17]. This is further supported by our DFT calculations using a plate capacitor model that considers both electrostatic and geometric changes to the $g$-tensor. The calculated $g_{z}$ variations of FePc (Fig. 1(e)) for each applied DC electric field after relaxation predict a frequency shift ($\sim$2 MHz) (Fig. 1(f)) two orders of magnitude smaller than our experimental observations. This is consistent with previous studies of pentacene molecules on thick insulating layers using atomic-force microscopy, i.e. in the absence of a polarized tunneling current [25].
Figure 2: ESR frequency ($f_{0}$) shift of a single Ti atom as a function of the DC bias. (a) Experimentally measured magnetic moments of Ti by varying external magnetic fields at different $V_{\text{DC}}$. The red line indicates the weighted average value. ESR conditions: $V_{\text{RF}}=30$ mV, $B_{z}=560$ mT, $T=1.8$ K. (b) Experimentally measured $f_{0}$ at different $V_{\text{DC}}$. The tip height was fixed by referring to the tunneling parameters of $V_{\text{DC}}=100$ mV, $I_{\text{set}}=35$ pA. (c) DFT-calculated $g_{z}$ of a Ti atom as a function of the out-of-plane electric field. (d) Simulated $f_{0}$-shift as a function of $V_{\text{DC}}$ based on the $g_{z}$-modulation in (c).
@@
Our analysis also rules out DC bias-induced vertical displacements of the adsorbate as a significant contributor to the observed $f_{0}$ shift. By examining the tunneling $I$-$z$ curve of FePc at various $V_{\text{DC}}$ values, we calculated that maintaining a constant $f_{0}$ across our $V_{\text{DC}}$ range ($20$ to $200$ mV) would require a tip-adsorbate distance change of approximately 20 pm. This implies an assumed experimental displacement of $\sim$70 pm/(V/nm) (calculated with an estimated tip-sample distance of $\sim$0.6 nm). Detailed information is provided in the SM (Sec. 4). Such a substantial vertical displacement should be easily detected by STM [26]; yet it was not observed in our experiments. Further supporting this conclusion, our DFT calculations predict that the mechanical displacement of FePc adsorbed on the MgO surface is only 0.44 pm/(V/nm) when subjected to a DC electric field. This suggests that the DC bias has negligible influence on the vertical adsorbate position in our $V_{\text{DC}}$ range and therefore does not contribute to the $f_{0}$ shift.
@@
To confirm the generality of the $V_{\text{DC}}$ induced resonance frequency shift, we studied individual Ti atoms for comparison. Similarly, the measured magnetic moments of Ti show no evident dependence on the $V_{\text{DC}}$ with an average value of 1.031 $\mu_{\text{B}}$ (Fig. 2(a). However, the resonance frequency increases by $\sim$150 MHz as $V_{\text{DC}}$ varies from 20 to 200 mV at fixed tip height (Fig. 2(b)). This distinct behavior again rules out $g$-factor changes as the primary mechanism. Our DFT calculations show minimal $g_{z}$ variation with DC electric field along the $z$-direction (Fig. 2(c)), and predict a slight frequency decrease ($\sim$20 MHz within ±$\pm{600}$ mV) with more positive $V_{\text{DC}}$ (Fig. 2(d)), contrary to experiments. Furthermore, the observed $\sim$150 MHz shift would require a tip adsorbate displacement of 50 pm for compensation (see SM. Sec. 4) in the $V_{\text{DC}}$ range of $20$ to $200$ mV, which corresponds to $\sim$170 pm/(V/nm) and has not been experimentally observed. In contrast, DFT predicts only 0.096 pm/(V/nm) for the atomic displacement. These findings conclusively demonstrate that neither $g$-factor modulation nor adsorbate displacement can satisfactorily explain the observed frequency shifts.
@@
Exchange field Model– Given that changes of the $g$-factor and vertical displacement are insufficient to explain the significant resonance frequency shift, we propose that it is mediated by the DC bias-dependent exchange field. The magnitude of $B_{\text{exch}}$ changes the probability of virtual electrons (or holes) exchange between the magnetic tip and the impurity and, as mentioned, it is not a mechanism exclusively for STM but rather generally appears in DC transport in magnetic tunnel junctions [27, 28, 29, 30, 31]. The resonance frequency shift $\Delta f$ caused by the exchange field is
@@
We now discuss the parameters that determine the exchange field. Under conditions typically found in low-temperature ESR-STM measurements, $\vec{B}_{\text{exch}}$ is given by (see SM. Sec. 5 and Ref. [20] for a detailed derivation):
$g\mu_{\text{B}}\vec{B}_{\text{exch}}=\frac{1}{2\pi}\left[\gamma_{\mathrm{T}}P_{\mathrm{T}}\log\left|\frac{eV_{\text{DC}}-\varepsilon-U}{eV_{\text{DC}}-\varepsilon}\right|\right]\vec{n}_{T}.$ (3)
[…]
@@
The exchange field depends exponentially on the tip-adsorbate distance through the coupling parameter $\gamma_{\text{T}}$, while the spin-polarization of the ferromagnetic tip $P_{\text{T}}$ affects the exchange field linearly. $P_{\text{T}}=0$ results in a vanishing exchange field, which underscores the importance of a polarized electrode for the $V_{\text{DC}}$-induced frequency shift  [^6] We note that the exchange field does not depend on the coupling to the substrate, $\gamma_{\text{S}}$, as the substrate is unpolarized.. Since only the product of $\gamma_{\text{T}}P_{\text{T}}$ enters Eq. (3) we can not determine them individually from a single measurement and we set $\left|P_{\text{T}}\right|=0.8$ following [42]. The apparent ferromagnetic and antiferromagnetic couplings between tip and magnetic impurity are determined by the sign of $P_{\text{T}}\cos\theta$. As an example, if $V_{\text{DC}}>0$, $P_{\text{T}}<0$ and $\theta\in(0,180)$ degrees, Eq. (3) yields a positive exchange field and the resonance frequency shifts toward larger values corresponding to a ferromagnetic coupling, while swapping the sign of $P_{T}$ yields antiferromagnetic behavior.
@@
This can be experimentally demonstrated by analyzing bistable tips with two opposing spin-polarizations. Bistable tips infrequently fluctuate between two magnetic states on a timescale that is slow compared to the signal integration of the ESR-STM [17, 43, 44]. Consequently, these tips produce spectra exhibiting two resonances that shift in opposite directions as a function of $V_{\text{DC}}$, as illustrated in Fig. 4(a). More importantly, both directions intersect at non-zero $V_{\mathrm{DC}}$, representing a symmetry point where electron and hole-like processes exactly cancel the exchange field (see Fig. 4(b) and SM, Fig. S7).
@@
Our model accounts for this phenomenon by considering both positive and negative spin-polarizations ±$\pm\left|P_{\text{T}}\right|$, yielding fits that accurately reproduce the experiment. This measurement directly reveals the crossing point, providing a better estimate of the $\varepsilon$ and $U$. Moreover, the calculated bare resonance frequency aligns well with this intersection, as shown by the red dashed line in Fig. 4(b). The crossing-point at finite bias also demonstrates the independence of the exchange field from the tunneling current, as a non-zero net exchange field can exist even at zero $V_{\text{DC}}$ and at zero current current. Crucially, the trend of opposing shifts could not be explained simply by a modulation of $g_{z}$ as the sign of the DC bias does not change when the polarization is inverted, enhancing the validity of our model.
@@
Figure 4: Resonance frequency shift of a Ti atom measured by bistable tip. (a) Experimentally measured (squares) and simulated (solid lines) frequency shift at varied $V_{\text{DC}}$ with a bistable tip showing two opposite signs of the spin-polarization (indicated by tip up, positive, and tip down, negative). In both (a) and (b), the red dash line represent the bare resonance frequency set by external magnetic field and $g$-factor. The blue curves in (b) shows the resonance frequencies calculated using Eqs. (2) and (3) with two opposing signs of the polarization (solid and dash lines). (c) Experimentally measured (diamonds and circles), simulated (blue and green solid lines) and bare (red and orange dashed lines) frequencies as a function of $V_{\text{DC}}$ with the external magnetic field along the out-of-plane ($B_{z}$) and in-plane ($B_{x}$) direction, respectively. Inset: geometry of the exchange field indicating the homodyne angle $\theta$ and the parallel (perpendicular) projections $B_{\text{exch}}^{\parallel}$ ($B_{\text{exch}}^{\perp}$) onto the spin $\vec{S}$. (d) Fitting results and other parameters used for the simulations in (a) and (c). ESR experimental parameters: (a) $V_{\text{RF}}=28$ mV, (b) $V_{\text{RF}}=30$ mV. The tip height at different $V_{\text{DC}}$ was fixed by referring to the tunneling condition of (a) $V_{\text{DC}}=100$ mV, $I_{\text{set}}=20$ pA, (c) $V_{\text{DC}}=100$ mV, $I_{\text{set}}=40$ pA.
@@
This implies that $g\mu_{B}B_{\text{exch}}^{\parallel}$ has no further dependence on the $g$-factor as the right hand side of Eq. (3) shows [28, 20]. Our data revealed different slopes in the shift of the resonance frequency of a Ti atom at varying $V_{\text{DC}}$ with out-of-plane and in-plane magnetic fields (Fig. 4(c)). We achieved good agreement with an angle of $\theta_{x}=32.71^{\circ}$, resulting in nearly identical adjustable parameters ($\gamma_{\text{T}}$, $\varepsilon$ and $U$) for both data sets, within the uncertainties of the fit (Fig. 4(d)). This consistency across different field orientations strongly supports the validity of our model, and contradicts the $g$-factor modulation or a piezoelectric displacement as the origin of the frequency shift, since they do not possess any dependence on the magnetic field orientation as they both depend on out-of-plane electric field modulation alone.
@@
A list of the best-fit parameters used for the model is given in Fig. 4(d). The impurity parameters $\varepsilon$ and $U$ are consistent within the fitting uncertainties, with detailed analysis of fit quality and uncertainty range of the fitting parameters provided in SM (Sec. 5, Fig. S8). Analysis of the four datasets shown in Fig. 4 yields average values of $\varepsilon=-237\pm 14$ meV and $U=568\pm 25$ meV. This is further confirmed by a statistical analysis of different datasets measured on Ti (see  SM. Figs. S3 and S4), indicating that the exchange mechanism is highly reproducible with different tips. Accurate determination of these parameters requires measurements in the strongly non-linear regime of the resonance frequency shift, which may be challenging to access due to the low activation barriers for adsorbate hopping and diffusion.
@@
Figure 5: Comparing the frequency shift of FePc and Ti with the same tip. (a) and (b) depict the experimental (circles and triangles) and simulated (solid lines) resonance frequencies as a function of $V_{\text{DC}}$ with the same tip for a FePc molecule and a Ti atom, respectively. The experimental values are extracted from Figs. 1 and 2. (c) Fitting parameters used for the FePc and Ti in (a) and (b), respectively. Since the tip here is different from the one of Fig. 4, $\theta=10.90^{\circ}$ gave the best fit of the ESR signal in terms of width and asymmetry contribution for both magnetic impurities.

### v3

@@ Abstract
Recent advances in scanning tunneling microscopy have enabled quantum-coherent control of single surface spins via all-electric electron spin resonance (ESR). Such control requires magnetoelectric coupling, since spin resonance is a magnetic effect. We show that a magnetic tip induces a bias-dependent exchange field on a localized Anderson impurity via virtual particle exchange with the magnetic lead. This field differs from Heisenberg exchange and can be tuned, reversed, or suppressed by the bias voltage. Our model reproduces bias-controlled resonance shifts for $S=1/2$ titanium atoms and Fe(II) phthalocyanine, enabling spin control via the exchange field and revealing the magnetoelectric mechanism behind all-electric ESR for spin-based quantum technologies.
@@
Introduction– The manipulation and control of individual quantum spins on surfaces is critical for advancing quantum information processing and spintronic devices at the nanoscale [1, 2]. While oscillating magnetic fields delivered via strip lines have traditionally been used to control the quantum states of spins, electric fields offer significant advantages, including faster switching speeds and more localized application through local gate electrodes [3, 4, 5, 6, 7, 8, 9]. To achieve this, it is crucial to understand how local electric fields can be converted to effective magnetic fields which can act on the spin state [10]. In this context, the recently developed combination of electron spin resonance (ESR) and scanning tunneling microscopy (STM) provides an excellent tool to probe and manipulate single atomic and molecular spins with atomic scale precision using purely AC electric fields [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Despite the success of this experimental technique in studying the interaction between electric fields and the local magnetic moments of spins on surfaces, open questions about the exact mechanism leading to the magnetoelectric coupling mechanism in all-electric ESR remain [41].
In this Letter, we focus on the bias-voltage dependent exchange interaction between a single electron spin localized in a magnetic adsorbate on a substrate and the magnetic tip of the STM. When the magnetic tip is sufficiently close to the impurity, the interaction between the spin and the tip is dominated by an exchange interaction, which leads to a shift in the resonance frequency when changing the tip-atom distance [20, 19, 22, 42, 36]. To understand the origin of this exchange interaction, we focus on its peculiar bias dependence at fixed tip-atom distance. In this situation, varying the DC bias yields a nonlinear resonance frequency shift, an inversion of the shift direction when the tip polarization changes sign, and a dependence of the shift on the angle between tip polarization and adsorbate spin. We also observe a sign reversal of the frequency shift when crossing a vanishing-shift point at non-zero DC bias. Originally, this bias-dependent resonance frequency shift was mainly attributed to $g$-factor modulation [42, 30, 43, 44], as the exchange field was expected to remain constant at fixed tip–atom distance.
We propose instead that the bias dependence originates entirely from the exchange field $B_{\mathrm{exch}}$, due to modulation of virtual particle exchange between the spin-polarized tip and the impurity spin. Our model is distinctively different from established Heisenberg-like exchange and based on a single-orbital Anderson impurity coupled to magnetic leads, which, under time-dependent modulation of the exchange field by an AC bias voltage, has been suggested as the origin of the all-electrical ESR mechanism in STM [45]. Here we will show that the component of the exchange-field parallel to the total magnetic field acting on the impurity spin is responsible for the frequency shift, which is related to spin-dependent energy renormalization observed in DC-transport in molecular Kondo quantum dots [46, 47, 48]. The good agreement between theory and experiment can alleviate the remaining doubts about the magnetoelectric coupling mechanism that drives ESR in the STM, pointing towards the exchange field as the most likely source.
@@
Experimental Results– In our experiment we utilized two adsorbates, FePc molecules and Ti atoms (Fig. 1(a)), which have electron spin $S=1/2$ when adsorbed on a bilayer of MgO grown on top of an Ag(100) substrate [28, 20]. The external magnetic field in this case was applied along out-of-plane direction ($B_{\mathrm{ext}}=B_{z}$) which, for a $S=1/2$, defines the spin quantization axis since the external magnetic field is much larger than the exchange field [45, 49]. Consequently, the resonance frequency is determined by the sum of external field and the exchange field component parallel to the quantization axis,
$hf_{0}=g\mu_{\mathrm{B}}(B_{\text{ext}}+B_{\text{exch}}^{\parallel}),$ (1)
@@
with $h$ the Planck constant, $g$ the $g$-factor and $\mu_{\mathrm{B}}$ the Bohr magneton. The $g$-factor can be determined by the frequency response to the total magnetic field [50]. We note that the $g$-factor depends on the adsorbed atomic or molecular species and the local crystal field environment and can deviate significantly from the free electron value. We performed ESR measurements on individual spins by adding a radio-frequency AC voltage ($V_{\text{RF}}$) in addition to $V_{\text{DC}}$ to the STM tip, as illustrated in Fig. 1(b). We follow the convention that $V_{\text{DC}}<0$ corresponds to probing the filled states of the sample.
@@
Figure 1: ESR frequency ($f_{0}$) shift of an individual FePc molecule and a Ti atom as a function of the DC bias. (a) STM image of the Fe, Ti atoms and the FePc molecule co-adsorbed on the MgO/Ag(100) surface. Scanning parameters: $V_{\text{DC}}=200$ mV, $I_{\text{set}}=20$ pA. (b) Schematic diagram of the ESR set-up at the tunneling junction, integrated into a plate capacitor model. $\vec{P}_{\mathrm{T}}$ represents the spin-polarization of the tip, forming an angle $\theta$ with the out-of-plane direction. (c) Experimentally measured magnetic moments of FePc at different $V_{\text{DC}}$. The red solid lines give the weighted average values of magnetic moments. (d) Corresponding resonance frequencies $f_{0}$ of FePc. The red dashed line indicates resonance frequencies calculated from DFT. (e) and (f) show the same data for Ti. The tip height was fixed with respect to the tunneling parameters of $V_{\text{DC}}=100$ mV, $I_{\text{set}}=20$ pA for FePc and $I_{\text{set}}=35$ pA for Ti. ESR conditions: $V_{\text{RF}}=20$ mV for FePc and $V_{\text{RF}}=30$ mV for Ti, $B_{z}=560$ mT, $T=1.8$ K.
@@
One possible explanation of the aforementioned $V_{\mathrm{DC}}$ dependence of the resonance frequency is a bias dependence of $g$ in Eq. (1), i.e. $g=g(V_{\mathrm{DC}})$ [30, 44]. To estimate the $g$-factor variation in FePc, we measured the magnetic moment $\mu$ under different external fields, as deviations of $\mu$ from $1\mu_{\mathrm{B}}$ imply a change of $g$ in Eq. (1). For FePc $\mu\approx 1\mu_{\mathrm{B}}$ [28, 51] and remains largely constant across different $V_{\text{DC}}$ values (Fig. 1(c)) indicating that the $g$-factor does not significantly change within the bias range. Simultaneously, we observed a significant monotonic decrease in resonance frequency $f_{0}$ with increasing $|V_{\text{DC}}|$ in the same conditions (Fig. 1(d)), ruling out a change of $g$-factor as dominant source of the frequency shift  [^1] We note that a precise determination of the $g$-factor from ESR-STM measurements is challenging due to competing influence of external and tip fields [42]. and calling for an alternative explanation. The lack of a strong electric-field dependence of the $g$-factor is further supported by our density functional theory (DFT) calculations using a plate capacitor model that considers both electrostatic and geometric changes to the $g$-tensor [53]. DFT predicts a negligible frequency shift ($\sim$2 MHz, red dashed lines in Fig. 1 (d)) two orders of magnitude smaller than our experimental observations. This is consistent with previous studies of pentacene molecules on thick insulating layers using atomic-force microscopy, i.e. in the absence of a polarized tunneling current [54].
@@
Further analysis also rules out DC bias-induced vertical displacements of the adsorbate as a significant contributor to the observed $f_{0}$ shift. By examining the tunneling II-$z$ curve of FePc at various $V_{\text{DC}}$ values, we calculated that maintaining a constant $f_{0}$ across our $V_{\text{DC}}$ range ($20$ to $200$ mV) would require a tip-adsorbate distance change of approximately 20 pm. This implies an assumed experimental displacement of $\sim$70 pm/(V/nm) (calculated with an estimated tip-sample distance of $\sim$0.6 nm [49]) Such a substantial vertical displacement should be easily detected by STM [55]; yet it was not observed in our experiments. Further supporting this conclusion, our DFT calculations predict that the mechanical displacement of FePc adsorbed on the MgO surface is only 0.44 pm/(V/nm) when subjected to a DC electric field. This suggests that the DC bias has negligible influence on the vertical adsorbate position in our $V_{\text{DC}}$ range and therefore does not contribute to the $f_{0}$ shift.
@@
To confirm the generality of the $V_{\text{DC}}$ induced resonance frequency shift, we studied individual Ti atoms for comparison. Similarly, the measured magnetic moments of Ti ($\mu_{Ti}$) show no evident dependence on the $V_{\text{DC}}$ with an average value of 1.031 $\mu_{\text{B}}$ (Fig. 1(e)), in general agreement with previous determinations of the $g$-factor [56]. However, the resonance frequency increases by $\sim$150 MHz as $V_{\text{DC}}$ varies from 20 to 200 mV at fixed tip height (Fig. 1(f)). This distinct behavior again rules out $g$-factor changes as the primary mechanism. DFT calculations performed in accordance with the calculations for FePc (red dashed line in Fig. 1(f)) show minimal $g_{z}$ variation with the DC electric field, and predict a slight frequency decrease ($\sim$20 MHz within $\pm{600}$ mV) with more positive $V_{\text{DC}}$, contrary to experiments. Furthermore, the observed $\sim$150 MHz shift would require a tip adsorbate displacement of 50 pm for compensation [49] in the $V_{\text{DC}}$ range of $20$ to $200$ mV, which corresponds to $\sim$170 pm/(V/nm) and has not been experimentally observed. In contrast, DFT predicts only 0.096 pm/(V/nm) for the atomic displacement. These findings conclusively demonstrate that neither $g$-factor modulation nor adsorbate displacement can satisfactorily explain the observed frequency shifts.
@@
Exchange field Model– Given that changes of the $g$-factor and vertical displacement are insufficient to explain the significant resonance frequency shift, we propose that it is mediated by a DC bias-dependent exchange field. The magnitude of $B_{\text{exch}}$ depends on the probability of virtual electrons (or holes) exchange between the magnetic tip and the impurity and is not a mechanism exclusive to STM but generally appears in DC transport in magnetic tunneling junctions [57, 58, 59, 60, 61].
The resonance frequency shift $\delta f$ caused by the exchange field is
@@
We now discuss the parameters that determine the exchange field. Under conditions typically found in low-temperature ESR-STM measurements, $\vec{B}_{\text{exch}}$ is given by [64, 45]:
$g\mu_{\text{B}}\vec{B}_{\text{exch}}=\frac{1}{2\pi}\left[\gamma_{\mathrm{T}}P_{\mathrm{T}}\ln\left|\frac{eV_{\text{DC}}-\varepsilon-U}{eV_{\text{DC}}-\varepsilon}\right|\right]\vec{n}_{T}.$ (3)
Equation (3) encapsulates all relevant transport parameters, primarily depending on the ionization ($\varepsilon$) and charge ($\varepsilon+U$) energies of the SAIM schematically depicted in Fig. 2(a). To correlate the charging energies of the adsorbate with the SAIM levels, we show a representative dII/dVV spectrum of an individual FePc molecule on MgO, and compare it with the simulated density of states, Fig. 2(b)  [^3] This interpretation relies on a change in the impurity’s charge state, $N\leftrightarrow N\pm 1$ [85].. The energies $\varepsilon$ and $\varepsilon+U$ determine the onset of the non-linear region of the resonance frequency shift, since $\vec{B}_{\text{exch}}$ increases logarithmically when the $eV_{\mathrm{DC}}$ approaches $\varepsilon$ or $\varepsilon+U$ (see Eq. (3) and Fig. 3(b)). Crucially, Eq. (3) shows that $\vec{B}_{\text{exch}}$ is independent from the resonance frequency, i.e from the Zeeman splitting, which is negligible compared to $\varepsilon$ and $\varepsilon+U$ for the standard values of the magnetic field used here (cf. Refs. [45, 58]), consistent with previous experimental findings [42, 30] and in contradiction with a frequency shift mediated by the $g$-factor [66]. Finally, we emphasize that the exchange field here is distinctively different from localized tip-spin models used to describe the interaction between tip and surface spin in previous works [67, 20] and provides a more natural description of the $V_{\mathrm{DC}}$-dependence. From the model point of view, it is the charge fluctuations of the Anderson model that provide this exchange field, which is absent in models that only describe spin fluctuations [68].
Figure 2: Energy diagram and dII/dVV spectrum for SAIM. (a) Energy-level diagram illustrating the key transport parameters governing the exchange field. The spin impurity splits into two states, $\downarrow$ (ground state) and $\uparrow$ (excited state), separated by the Zeeman energy. Filled and empty arrows denote singly occupied electron states and hole states, respectively. The applied DC bias is $eV_{\text{DC}}=\mu_{S}-\mu_{T}$. (b) Simulated density of states (DOS, left) and experimental dII/dVV spectrum (right) of a FePc molecule on MgO.
@@
The exchange field depends exponentially on the tip-adsorbate distance through the coupling $\gamma_{\text{T}}$, while the spin-polarization of the ferromagnetic tip, $P_{\text{T}}$, affects the exchange field linearly. $P_{\text{T}}=0$ results in a vanishing exchange field, which underscores the importance of a polarized electrode for the $V_{\text{DC}}$-induced frequency shift  [^6] We note that the exchange field does not depend on the coupling to the substrate, $\gamma_{\text{S}}$, as it is unpolarized.. Since only the product of $\gamma_{\text{T}}P_{\text{T}}$ enters Eq. (3) we cannot determine them individually from a single measurement and we set $\left|P_{\text{T}}\right|=0.8$ following [74]. The apparent ferromagnetic or antiferromagnetic coupling between the tip and the magnetic impurity is determined by the sign of $P_{\text{T}}\cos\theta$. For example, when $V_{\text{DC}}>0$, $P_{\text{T}}<0$, and $\theta\in(0,180^{\circ})$, Eq. (3) yields a positive exchange field, and the resonance frequency shifts toward larger values, a ferromagnetic behavior. Conversely, reversing the sign of $P_{\text{T}}$ leads to an antiferromagnetic coupling.
@@
This can be experimentally demonstrated by analyzing bistable tips with two opposing spin-polarizations. Bistable tips infrequently fluctuate between two magnetic states on a timescale that is slow compared to the signal integration of the ESR-STM [42, 75, 76]. Consequently, these tips produce spectra exhibiting two resonances that shift in opposite directions as a function of $V_{\text{DC}}$, as illustrated in Fig. 3(a). More importantly, both directions intersect at non-zero $V_{\mathrm{DC}}$, representing a symmetry point where electron and hole-like processes exactly cancel the exchange field [64].
@@
Our model accounts for this phenomenon by considering both positive and negative spin-polarizations $\pm\left|P_{\text{T}}\right|$, yielding fits that accurately reproduce the experiment. This measurement directly reveals the crossing point, providing a better estimate of $\varepsilon$ and $U$. Moreover, the calculated bare resonance frequency aligns well with this intersection, as shown by the red dashed line in Fig. 3(b). The crossing-point at finite bias also demonstrates the independence of the exchange field from the tunneling current, as a non-zero net exchange field can exist even at zero $V_{\text{DC}}$ and at zero current current. Crucially, the trend of opposing shifts could not be explained simply by a modulation of $g_{z}$ as the sign of the DC bias does not change when the polarization is inverted, enhancing the validity of our model.
@@
Figure 3: Resonance frequency shift of a Ti atom measured by bistable tip. (a) Experimentally measured (squares) and simulated (solid lines) frequency shift at varied $V_{\text{DC}}$ with a bistable tip showing two opposite signs of the spin-polarization (tip up, positive, and tip down, negative). In both (a) and (b), the red dashed line represent the bare resonance frequency set by external magnetic field and $g$-factor. The blue curves in (b) shows the resonance frequencies calculated using Eqs. (2) and (3) with two opposing signs of the polarization (solid and dashed lines). (c) Experimentally measured (diamonds and circles), simulated (blue and green solid lines) and bare (red and orange dashed lines) frequencies vs $V_{\text{DC}}$ with the external magnetic field along the out-of-plane ($B_{z}$) and in-plane ($B_{x}$) direction, respectively. Inset: geometry of the exchange field indicating the homodyne angle $\theta$ and the parallel (perpendicular) projections $B_{\text{exch}}^{\parallel}$ ($B_{\text{exch}}^{\perp}$) onto the spin $\vec{S}$. (d) Parameters used for the simulations in (a) and (c). ESR experimental parameters: (a) $V_{\text{RF}}=28$ mV, (b) $V_{\text{RF}}=30$ mV. The tip height at different $V_{\text{DC}}$ was fixed by referring to the tunneling condition of (a) $V_{\text{DC}}=100$ mV, $I_{\text{set}}=20$ pA, (c) $V_{\text{DC}}=100$ mV, $I_{\text{set}}=40$ pA.
@@
[…]
@@
As displayed by the list of the best-fit parameters used for the model in Fig. 3(d), the impurity parameters $\varepsilon$ and $U$ are consistent within the fitting uncertainties [78]. Accurate determination of these parameters requires measurements in the strongly non-linear regime of the resonance frequency shift, which may be challenging to access due to the low activation barriers for adsorbate hopping and diffusion.
@@
Figure 4: Comparing the frequency shift of FePc and Ti measured with the same tip. (a) and (b) depict the experimental (circles and triangles) and simulated (solid lines) resonance frequencies as a function of $V_{\text{DC}}$ with the same tip for a FePc molecule and a Ti atom, respectively. The experimental values are extracted from Figs. 1 and 2. (c) Parameters used in the transport simulations for FePc and Ti in (a) and (b), respectively. Since the tip used here differs from that in Fig. 3, a different angle $\theta$ is chosen, approximately one third of $\theta_{x}$, which provides a reasonable fit of the ESR signal in terms of both linewidth and asymmetry.